# Beyond Trial-and-Error: Predicting User Abandonment After a Moderation Intervention

### v2

@@ III-B Task specification
In this work, we aim to predict those users of banned subreddits who abandoned the platform —that is, who ceased all activity on Reddit— after the Great Ban. Given the novelty of the problem and for the sake of simplicity, we perform two variants of the classification task $C$ imposing $m=n$ so that $P^{-}$ and $P^{+}$ have the same duration, but with different abandonment time windows $P^{*}$. We call these tasks $C^{H}$ for hard abandonment and $C^{S}$ for soft abandonment, which are defined as:
$C^{H}$ :=$:=\{C|P^{*}=P^{+}\}$
$C^{S}$ :=$:=\{C|P^{*}\subset{}P^{+}\;\text{and}\;t_{+a}\gg{}t_{0}\}.$
@@
Based on previous research that analyzed the medium-term effect of moderation interventions on Reddit [13, 11], we determined the duration of both $P^{-}$ and $P^{+}$ to be 7 months (210 days), for a total time frame of 421 days (including the day $t_{0}$ of the Great Ban). Hence, for $C^{H}$ the duration of $P^{*}$ is also 7 months. On the other hand, for $C^{S}$ we selected an offset $t_{+a}$ of 4 months, with the duration of the corresponding $P^{*}$ being 3 months (92 days), based on a precursory analysis of different offsets and class balances between non-abandoning and abandoning users.
@@
TABLE I: List of the 15 banned subreddits used for the analysis, sorted by number of active users. Data from these subreddits constitutes dataset DB-B.
IV Data Collection and Preparation
For this study we use the dataset from [11], consisting of 16M Reddit comments made by 16,828 different users that were affected by the Great Ban.
IV-A Banned subreddits selection
Each user in the dataset was an active participant before the ban in at least one of 15 selected subreddits. The selection of the 15 subreddits was done as follows. Out of the 2,000 subreddits shut during the Great Ban, Reddit only publicly disclosed the names of the 10 largest ones. [^1] https://www.redditstatic.com/banned-subreddits-june-2020.txt (accessed: 03/15/2024) For all remaining banned subreddits, Reddit only provided a list of partially obfuscated names. The authors of [19] deciphered the list, removed all private subreddits as well as all those having less than 2,000 active users. Only 5 subreddits remained from the obfuscated list, which [19] and [11] studied in addition to the 10 publicly disclosed ones. Table I provides the list of the 15 subreddits on which our study is focused, together with some descriptive statistics.
@@ IV-A Banned subreddits selection
IV-B Active users selection
For each of the 15 selected subreddits, the authors of [11] gathered all comments posted therein between December 2019 and June 2020, which resulted in a total of 8M comments posted by approximately 194K distinct users. As shown in Figure 2, the timeframe used for the data collection spans 7 months leading up to the Great Ban, providing a strong reference for the activity levels of the affected users before the moderation intervention [14]. The initial set of 194K users was later filtered so as to retain only those users who showed consistent activity in at least one of the 15 selected subreddits. In detail, only those users who posted at least one comment each month were retained. In [11], this filtering step was useful to obtain meaningful post-hoc estimations of the effect of the Great Ban on the activity of the users. Since here we also study changes in user activity, albeit in a predictive fashion, this filtering step is suitable for our analysis as well. In addition, bots were also removed from the dataset. First, we followed reference practices in literature and we removed all accounts that posted two or more comments in less than a second [29]. Subsequently, we manually analyzed a random sample of 1,000 removed accounts to ensure that only bots were removed from the dataset. A similar analysis also revealed that among the remaining users in the dataset there were no clearly distinguishable bots. As a result of these filtering and validation steps, we obtained the dataset DB-B composed of 2.2M comments posted within the 15 selected subreddits by 16,828 distinct users, as summarized in Table I.
IV-C Data from outside the banned subreddits
Evaluating (or predicting) the effect of a moderation intervention involves comparing data from before and after the intervention itself. This requirement also surfaces from the definitions that we gave in Section III. However, no activity exists post-intervention within the banned subreddits, since they were all permanently shut. Therefore, data about the behavior of the affected users outside of the banned subreddits must be used. For this reason, the authors in [11] collected all comments made by the 16,828 selected users outside of the 15 banned subreddits over a wide timeframe spanning 7 months before and 7 months after the Great Ban, as depicted in Figure 2. This additional data collection yielded approximately 13.8M comments, of which 8.2M (59%) were posted before the ban and constitute dataset DN-B while 5.6M (41%) were posted afterwards and constitute dataset DN-A. We used the data from before the ban (DN-B) to compute our machine learning features, while the data posted after the ban (DN-A) allows assigning ground-truth labels to the users based on their post-ban activity. Notably, out of the 16,828 initially selected users, 288 (1.7%) did not have any activity outside of the 15 banned subreddits and were thus discarded.
V Machine Learning Approach
@@
The formulation of hard and soft abandonment prediction given in Section III-B cast both problems as binary classification tasks. Our approach to these tasks is sketched in Figure 3 and summarized in the remainder of this section, while the implementation details and the experimental settings are described in Section VII.
@@
We begin by extracting user-level features from the portion of the dataset related to the activity of the selected users before the ban, both within (DB-B) and outside (DN-B) of the banned subreddits. Before further processing this data we perform outliers removal, as this helps enhancing the robustness and generalization of the trained models by mitigating the potential distortions and biases caused by extreme data points [30]. Data about user activities outside of the banned subreddits after the ban (DN-A) is used to assign ground-truth labels to the users. User-level features are then merged with the ground-truth labels and the resulting dataset is split into a training and a test set. The test set is only used for the final evaluation of the optimized models. Instead, the training set is used for feature selection and for optimizing and training the models. Given that we experiment with multiple classification models, each with its own characteristics and hyperparameters, the feature selection and optimization steps of our machine learning pipeline ensure that each model operates in optimal conditions. In other words, our approach ensures meaningful and fair comparisons between the different models. The first step in the model optimization process involves selecting an adequate number and set of features for each model. Specifically, we rank all features and we select the top-N ones to reduce the dimensionality of our models and to eliminate redundant and correlated features. For each model we experiment with multiple numbers N of selected features, as different models might benefit from a smaller or larger number of features. To have accurate and robust estimates of the model performances with the different sets of features, we carry out a 10-fold cross-validation on the training set. At each iteration, we rescale features so that they all cover a similar range of values. Rescaling ensures that no feature dominates the learning process due to overly large magnitude, and also contributes to speeding up training times  [31, 32]. At each iteration we also rebalance the training set. In fact, both the hard and soft abandonment tasks are heavily imbalanced, given that abandoning users constituted a small minority of all users (i.e., 14.9% for hard abandonment and 26.9% for soft abandonment). We note that this imbalance is common for tasks concerning the prediction of human behavior [33]. Therefore, by rebalancing the training set we reduce the bias that the trained models might exhibit towards those users who did not abandon the platform (i.e., the majority class). The averaged results of the 10-fold cross-validation process are used to select the optimal number and set of features for each model. Up to now, all models were trained and evaluated with default hyperparameters. In the last step of model optimization, for each model we perform a grid search with cross-validation over its hyperparameters so as to choose the best combination of them. Finally, we use the selected set of features and hyperparameters to train an optimized version of each model on the whole training set. Feature scaling and rebalancing are applied also for hyperparameter tuning and model training. Each optimized model is then evaluated on the held-out test set. Evaluation results and comparisons are presented in Section VIII.
VI Feature Engineering
@@
VI-A Activity features
Features in this class describe the overall level of engagement and participation of a user within Reddit. The motivation for including activity features is straightforward, given our goal of predicting user abandonment—that is, the interruption of posting activity on the platform. Moreover, previous descriptive studies on the effects of moderation interventions, such as those reviewed in Section II-A, showed that bans frequently cause a reduction in user activity [11, 19, 13], which supports the adoption of this type of features in our task. Some examples of the activity features that we implemented are the total number of comments posted by a user, the average time between the sharing of two subsequent comments, the slope of the trend of posted comments in either the banned and the non-banned subreddits, which captures whether user activity in certain subreddits was increasing or decreasing prior to the ban. Other features of this class capture possible peculiar characteristics of the comments, one of them being the number of stickied comments. This feature was included as it was among the strongest predictors of ban evasion in a previous study [28].
VI-B Toxicity features
@@
VI-C Writing style features
Several previous works demonstrated that certain users exhibit linguistic changes on a platform after suffering a moderation intervention. For example, [6] noted changes in the use of the first and third plural pronouns, as well as in word choices, after a ban. Similarly, [19] found that users who remained on Reddit after a ban exhibited reduced use of in-group language. Based on these preliminary descriptive results, here we assess the extent to which basic linguistic and writing style features provide predictive information about the activity of the users after a ban. Among the features that we computed are counts of the different parts-of-speech used and readability scores such as the well-known Flesch-Kincaid Grade Level and the SMOG Index. Each of these scores were computed for each user comment and subsequently aggregated at the user level by computing the mean, minimum, maximum, and standard deviation.
VI-D Relational features
@@
VII Experiments and Settings
VII-A Comparisons
We experiment with multiple reference classification models to solve the hard and soft abandonment tasks, including Naive Bayes (NB), K-Nearest Neighbors (KNN), Decision Tree (DT), Random Forest (RF), Adaptive Boosting (AB), Gradient Boosting (GB), and Support Vector Machine (SVM). This selection spans the main families of classification algorithms, encompassing instance-based, probabilistic, ensemble-based, and discriminative methods of different complexities, thus providing an extensive exploration of many learning paradigms. Furthermore, similar choices of classification models were made in recent works on related tasks, such as the detection of different forms of online harms [42, 43, 44, 45]. In addition to the previous classification models, we also implement three simple baselines for further comparison. The Stratified baseline generates predictions by following the ground-truth label distribution, without taking any feature into account. The DT Ratio baseline employs a decision tree that receives a single feature in input representing the ratio of comments made by a user in the non-banned subreddits over the comments made that user in the banned ones. The rationale for this baseline is that users who are more active on non-banned subreddits are less prone to abandon the platform. Finally, The DT Trend baseline employs a decision tree that receives a single feature in input representing the trend of the number of monthly comments by a user before the ban. The reason for including this baseline is that users with a decreasing posting trend are more likely to abandon the platform. For the sake of simplicity, we did not apply any preprocessing steps to the baselines. Finally, we remark that we resort to using these baselines instead of more powerful approaches because the novelty of our task makes it so that no previous work exists on predicting the effects of a moderation intervention.
VII-B Machine learning pipeline
@@ VII-B1 Outliers removal
VII-B2 Train-test split
We use a standard 80/20 split of our dataset into the training and test sets, following the Pareto principle. Data in the two sets is stratified according to the class labels.
VII-B3 Cross-validation
@@
VII-B4 Feature scaling and rebalancing
We perform feature scaling and data rebalancing during feature selection, hyperparameter tuning, and model training. Feature scaling is applied first, via Z-score standardization, so that each feature has mean $\mu=0$ and standard deviation $\sigma=1$. Rebalancing is applied last, as suggested in previous literature [47]. For rebalancing, we leverage both oversampling and undersampling techniques to mitigate the large class imbalance of our dataset. We perform oversampling with SMOTE, a state-of-the-art technique that creates new artificial samples of the minority class [47]. Instead, we use Random Under Sampling to randomly discard samples from the majority class [48]. For both tasks we use the two techniques in combination to reach a final fixed imbalance of 60/40 between the majority and minority class, down from 85/15 for hard abandonment and 73/27 for soft abandonment. This approach allows us to reduce the skewness of the class distribution without creating too many artificial samples that could introduce bias, and without discarding too much data that could result in degraded model performance.
VII-B5 Feature selection
We perform feature selection via ANOVA F-value, a technique that was shown to achieve good performance in recent works on the detection of malicious online user activities [49]. The ANOVA tests are used to obtain an F-value for each feature, expressing the likelihood for the feature to be predictive for the task at hand. Then, features are ranked based on their respective F-value and the top-k are selected. Out of 142 total features, for each trained model we vary k from 10 to 80 with a step of 10, and evaluate the performance of the resulting models as part of the 10-fold cross-validation. At the end of the 10-fold cross-validation process, for each model we select the number and set of features that maximizes its F1 score on the positive class, so that for each model, only its best set of features is used in the following steps.
VII-B6 Hyperparameter tuning

### v3

@@ 3.2 Task specification
In this work, we aim to predict those users of banned subreddits who abandoned the platform —that is, who ceased activity on Reddit— after the Great Ban. Given the novelty of the problem and for the sake of simplicity, we perform two variants of the classification task $C$. In both variants we set $m=n$ so that $P^{-}$ and $P^{+}$ have the same duration, but we use two different settings for $t_{a}$, which results in two different abandonment periods $P^{a}$:
$C^{H}$ :=$:=\{C\;|\;t_{+a}=t_{0}\;\implies\;P^{a}=P^{+}\}\quad\text{({hard abandonment})}$
$C^{S}$ :=$:=\{C\;|\;t_{+a}\gg t_{0}\;\implies\;P^{a}\subset P^{+}\}\quad\text{({soft abandonment})}.$
The hard abandonment task $C^{H}$ follows a strict definition, where a user is considered non-abandoning if they exhibit any activity on the platform after the intervention. Conversely, the soft abandonment task $C^{S}$ introduces a more flexible criterion. In this case, a user may continue posting for a limited time after the intervention —up to $t_{+a}$— and still be classified as abandoning if they become inactive beyond this point. This approach allows for a more nuanced distinction, recognizing that some users may gradually disengage rather than leaving the platform immediately.
@@
Based on previous research that analyzed the medium-term effects of moderation interventions on Reddit [13, 11], we determined the duration of $P^{-}$ and $P^{+}$ to be 210 days, for a total time frame of 421 days (including the day $t_{0}$ of the Great Ban). The selected time frame provides a convenient and practical period that can be divided into months (7 months) or weeks (30 weeks). Hence, for $C^{H}$ the duration of $P^{a}=P^{+}=$ 7 months. On the other hand, for $C^{S}$ we selected the offset $t_{+a}=$ 4 months, with the duration of the corresponding abandonment period $P^{a}=$ 3 months (92 days), based on a precursory analysis of different offsets and class balances between non-abandoning and abandoning users. The 4-month inactivity threshold balances a long enough time to indicate meaningful disengagement while avoiding possible short-term fluctuations in user activity.
@@
Table 1: List of the 15 banned subreddits used for the analysis, sorted by number of active users. Data from these subreddits constitutes dataset DB-B.
4 Data Collection and Preparation
For this study we use the dataset from [11, 18], consisting of 16 million Reddit comments made by 16,828 different users that were affected by the Great Ban.
4.1 Banned subreddits selection
Each user in the dataset was an active participant before the ban in at least one of 15 selected subreddits. The selection of the 15 subreddits was done as follows. Out of the 2,000 subreddits shut during the Great Ban, Reddit only publicly disclosed the names of the 10 largest ones. [^1] https://www.redditstatic.com/banned-subreddits-june-2020.txt (accessed: 03/15/2024) For all remaining banned subreddits, Reddit only provided a list of partially obfuscated names. M. Trujillo et al. [21] deciphered the list, removed all private subreddits as well as all those having less than 2,000 active users. Only 5 subreddits remained from the obfuscated list, which M. Trujillo et al. [21] and Cima et al. [11] studied in addition to the 10 publicly disclosed ones. Table 1 provides the list of the 15 subreddits on which our study is focused, together with some descriptive statistics.
@@ 4.1 Banned subreddits selection
4.2 Active users selection
For each of the 15 selected subreddits, Cima et al. [11] gathered all comments posted therein between December 2019 and June 2020, which resulted in a total of 8 million comments posted by approximately 194K distinct users. As shown in Figure 3, the timeframe used for the data collection spans 7 months leading up to the Great Ban, providing a strong reference for the activity levels of the affected users before the moderation intervention [14]. The initial set of 194K users was later filtered so as to retain only those users who showed consistent activity in at least one of the 15 selected subreddits. In detail, only those users who posted at least one comment each month were retained. In [11], this filtering step was useful to obtain meaningful post-hoc estimations of the effect of the Great Ban on the activity of the users. Since here we also study changes in user activity, albeit in a predictive fashion, this filtering step is suitable for our analysis as well. In addition, bots were also removed from the dataset. First, we followed reference practices in literature and we removed all accounts that posted two or more comments in less than a second [46]. Subsequently, we manually analyzed a random sample of 1,000 removed accounts to ensure that only bots were removed from the dataset. A similar analysis also revealed that among the remaining users in the dataset there were no clearly distinguishable bots. As a result of these filtering and validation steps, we obtained the dataset DB-B composed of 2.2 million comments posted within the 15 selected subreddits by 16,828 distinct users, as summarized in Table 1.
4.3 Data from outside the banned subreddits
Evaluating (or predicting) the effect of a moderation intervention involves comparing data from before and after the intervention itself. This requirement also surfaces from the definitions that we gave in Section 3. However, no activity exists post-intervention within the banned subreddits, since they were all permanently shut. Therefore, data about the behavior of the affected users outside of the banned subreddits must be used. For this reason, Cima et al. [11] collected all comments made by the 16,828 selected users outside of the 15 banned subreddits over a wide timeframe spanning 7 months before and 7 months after the Great Ban, as depicted in Figure 3. This additional data collection yielded approximately 13.8 million comments, of which 8.2 million (59%) were posted before the ban and constitute dataset DN-B while 5.6 million (41%) were posted afterwards and constitute dataset DN-A. We used the data from before the ban (DN-B) to compute our machine learning features, while the data posted after the ban (DN-A) allows assigning ground-truth labels to the users based on their post-ban activity. Notably, out of the 16,828 initially selected users, 288 (1.7%) did not have any activity outside of the 15 banned subreddits and were thus discarded.
@@ 4.3 Data from outside the banned subreddits
5 Machine Learning Approach
@@
The formulation of hard and soft abandonment prediction given in Section 3.2 cast both problems as binary classification tasks. Our approach to these tasks is sketched in Figure 4 and summarized in the remainder of this section, while the implementation details and the experimental settings are described in Section 7.
@@
We begin by extracting user-level features from the portion of the dataset related to the activity of the selected users before the ban, both within (DB-B) and outside (DN-B) of the banned subreddits. Before further processing this data we perform outliers removal, as this helps enhancing the robustness and generalization of the trained models by mitigating the potential distortions and biases caused by extreme data points [47]. Data about user activities outside of the banned subreddits after the ban (DN-A) is used to assign ground-truth labels to the users. User-level features are then merged with the ground-truth labels and the resulting dataset is split into a training and a test set. The test set is only used for the final evaluation of the optimized models. Instead, the training set is used for feature selection and for optimizing and training the models. Given that we experiment with multiple classification models, each with its own characteristics and hyperparameters, the feature selection and optimization steps of our machine learning pipeline ensure that each model operates in optimal conditions. In other words, our approach ensures meaningful and fair comparisons between the different models. The first step in the model optimization process involves selecting an adequate number and set of features for each model. Specifically, we rank all features and we select the top-N ones to reduce the dimensionality of our models and to eliminate redundant and correlated features. For each model we experiment with multiple numbers N of selected features, as different models might benefit from a smaller or larger number of features. To have accurate and robust estimates of the model performances with the different sets of features, we carry out a 10-fold cross-validation on the training set. At each iteration, we rescale features so that they all cover a similar range of values. Rescaling ensures that no feature dominates the learning process due to overly large magnitude, and also contributes to speeding up training times  [48, 49]. At each iteration we also rebalance the training set. In fact, both the hard and soft abandonment tasks are heavily imbalanced, given that abandoning users constituted a small minority of all users (i.e., 14.9% for hard abandonment and 26.9% for soft abandonment). We note that this imbalance is common for tasks concerning the prediction of human behavior [50]. Therefore, by rebalancing the training set we reduce the bias that the trained models might exhibit towards those users who did not abandon the platform (i.e., the majority class). The averaged results of the 10-fold cross-validation process are used to select the optimal number and set of features for each model. Up to now, all models were trained and evaluated with default hyperparameters. In the last step of model optimization, for each model we perform a grid search with cross-validation over its hyperparameters so as to choose the best combination of them. Finally, we use the selected set of features and hyperparameters to train an optimized version of each model on the whole training set. Feature scaling and rebalancing are applied also for hyperparameter tuning and model training. Each optimized model is then evaluated on the held-out test set. Evaluation results and comparisons are presented in Section 8.
6 Feature Engineering
@@
6.1 Activity features
Features in this class describe the overall level of engagement and participation of a user within Reddit. The motivation for including activity features is straightforward, given our goal of predicting user abandonment—that is, the interruption of posting activity on the platform. Moreover, previous descriptive studies on the effects of moderation interventions, such as those reviewed in Section 2.1, showed that bans frequently cause a reduction in user activity [11, 21, 13, 18], which supports the adoption of this type of features in our task. Some examples of the activity features that we implemented are the total number of comments posted by a user, the average time between the sharing of two subsequent comments, the slope of the trend of posted comments in either the banned and the non-banned subreddits, which captures whether user activity in certain subreddits was increasing or decreasing prior to the ban. Other features of this class capture possible peculiar characteristics of the comments, one of them being the number of stickied comments. This feature was included as it was among the strongest predictors of ban evasion in a previous study [31].
6.2 Toxicity features
@@
6.3 Writing style features
Several previous works demonstrated that users may exhibit linguistic changes after suffering a moderation intervention. For example, [7] noted changes in the use of the first and third plural pronouns, as well as in word choices, after a ban. Similarly, [21] found that users who remained on Reddit after a ban exhibited reduced use of in-group language. Other prior research has shown that moderation interventions can lead to linguistic shifts, including changes in word choices, sentiment, and rhetorical strategies. Studies have observed that moderated users may adjust their language to evade detection, adopt more aggressive or explicit speech in response to relaxed moderation policies, or even develop lexical variations of restricted terms to bypass enforcement mechanisms. These findings suggest that subtle linguistic cues may reflect user adaptation, disengagement, or defiance following an intervention [58, 21, 59]. Based on these preliminary descriptive results, here we assess the extent to which basic linguistic and writing style features provide predictive information about the activity of the users after a ban. Among the features that we computed are counts of the different parts-of-speech used and readability scores such as the well-known Flesch-Kincaid Grade Level and the SMOG Index. Each of these scores were computed for each user comment and subsequently aggregated at the user level by computing the mean, minimum, maximum, and standard deviation.
6.4 Relational features
@@
7 Experiments and Settings
7.1 Comparisons
We experiment with multiple reference classification models to solve the hard and soft abandonment tasks, including Naive Bayes (NB), K-Nearest Neighbors (KNN), Decision Tree (DT), Random Forest (RF), Adaptive Boosting (AB), Gradient Boosting (GB), and Support Vector Machine (SVM). This selection spans the main families of classification algorithms, encompassing instance-based, probabilistic, ensemble-based, and discriminative methods of different complexities, thus providing an extensive exploration of many learning paradigms. Furthermore, similar choices of classification models were made in recent works on related tasks, such as the detection of different forms of online harms [61, 62, 63, 64]. In addition to the previous classification models, we also implement five simple baselines for further comparison. The Stratified baseline generates predictions by following the ground-truth label distribution, without taking any feature into account. The DT Ratio baseline employs a decision tree that receives a single feature in input representing the ratio of comments made by a user in the non-banned subreddits over the comments made by that user in the banned ones. The rationale for this baseline is that users who are more active on non-banned subreddits are less prone to abandon the platform. The DT Trend baseline employs a decision tree that receives a single feature in input representing the trend of the number of monthly comments by a user before the ban. The reason for including this baseline is that users with a decreasing posting trend are more likely to abandon the platform. Finally, to test the effectiveness of our machine learning pipeline, we train a Naive Bayes (NB-NFE) and a Decision Tree (DT-NFE) model by only performing feature standardization, without any other step of our pipeline —namely, rebalancing, feature selection, and hyperparameter tuning. For the sake of simplicity, we do not apply any preprocessing steps to the baselines. Finally, we remark that we resort to using these baselines instead of more powerful approaches because the novelty of our task makes it so that no previous work exists on predicting the effects of a moderation intervention.
7.2 Machine learning pipeline
@@ 7.2.1 Outliers removal
7.2.2 Train-test split
We use a common 80/20 split of our dataset into the training and test sets, following the Pareto principle. Data in the two sets is stratified according to the class labels.
7.2.3 Cross-validation
@@
7.2.4 Feature scaling and rebalancing
During the feature selection, hyperparameter tuning, and model training steps, we perform feature scaling for all models. Additionally, for all models except for the ensemble-based ones —namely, RF, AB, and GB— we also perform data rebalancing. For ensemble-based models, removing data rebalancing often improves performance, as also confirmed by our experimental results, because these models naturally handle class imbalance and can be negatively affected by data rebalancing, which might distort data distributions [66]. Feature scaling is applied first, via Z-score standardization, so that each feature has mean $\mu=0$ and standard deviation $\sigma=1$. Rebalancing is applied last, as suggested in previous literature [67]. For rebalancing, we leverage both oversampling and undersampling techniques to mitigate the large class imbalance of our dataset. We perform oversampling with SMOTE, a state-of-the-art technique that creates new artificial samples of the minority class [67]. Instead, we use Random Undersampling to randomly discard samples from the majority class [68]. For both tasks we use the two techniques in combination to reach a final fixed imbalance of 60/40 between the majority and minority class, down from 85/15 for hard abandonment and 73/27 for soft abandonment. This approach allows us to reduce the skewness of the class distribution without creating too many artificial samples that could introduce bias, and without discarding too much data that could result in degraded model performance.
7.2.5 Feature selection
We perform feature selection for all models except for the ensemble-based ones. Indeed, the latter perform internal feature selection, reducing the need for explicit preprocessing, and are robust to redundant features, unlike simpler models [69]. Feature selection is achieved with ANOVA F-value, a technique that was shown to achieve good performance in recent works on the detection of malicious online user activities [70]. The ANOVA tests are used to obtain an F-value for each feature, expressing the likelihood for the feature to be predictive for the task at hand. Then, features are ranked based on their respective F-value and the top-k are selected. Out of 142 total features, for each trained model we vary k from 10 to 80 with a step of 10, and evaluate the performance of the resulting models as part of the 10-fold cross-validation. At the end of the 10-fold cross-validation process, for each model we select the number and set of features that maximizes its F1 score on the positive class, so that for each model, only its best set of features is used in the following steps.
7.2.6 Hyperparameter tuning